OpenStudy (james.b):

Evan spent 20 hours doing homework last week. This week he spent 25 hours doing homework. He says that he spent 125% more time doing homework this week. Is he correct? Show your work to justify your decision.

OpenStudy (anonymous):

Yes Evan is correct Let 20hrs=100/100 i.e 100% Then 25hrs=(100/100)+(25/100) i.e 125%